# A Model Predictive Control Scheme for Flight Scheduling and Energy Management of Electric Aviation Networks

### v1

@@ Abstract
This paper presents a Model Predictive Control (MPC) scheme for flight scheduling and energy management of electric aviation networks, where electric aircraft transport passengers between electrified airports equipped with sustainable energy sources and battery storage, with the goal of minimizing grid dependency. Specifically, we first model the aircraft flight and charge scheduling problem jointly with the airport energy management problem, explicitly accounting for local weather forecasts. Second, we frame the minimum-grid-energy operational problem as a mixed-integer linear program and solve it in a receding horizon fashion, where the route assignment and charging decisions of each aircraft can be dynamically reassigned to mitigate disruptions. We showcase the proposed MPC scheme on real-world data taken from conventional flights and weather conditions in the Dutch Leeward Antilles. Our results show that MPC can effectively guarantee operation of the network by efficiently re-assigning flights and rescheduling aircraft charging, whilst maximizing the efficiency of the on-site energy systems.
Index Terms:
Electric Aviation, Model Predictive Control
@@ I Introduction
In the following we often concatenate variables into column vectors denoted with bold letters as in $\mathbf{v}=[v(1)\;v(2)\;\dots\;v(n)]^{\top}$, such that the dimension of $\mathbf{v}$ would be $n\times 1$. A column vector of all ones with dimension $m$ is denoted $\mathbf{1}_{m}$, the $m\times m$ identity matrix is denoted $I_{m}$, and $\Lambda_{m}$ denotes an $m\times m$ lower triangular matrix of ones. Matrices are denoted with capital letters $M$, and the element in the $m$-th row and $n$-th column is indexed with $[M]_{mn}$. We use $\otimes$ for the Kronecker product operator.
@@ II-B Aviation Network Model
We extract all flights that start in the upcoming horizon from the given flight schedule and construct a set of flight edges that connect nodes of different airports. For every flight we allow a maximum possible flight delay of $\delta t$, an integer multiple of $\Delta t$, meaning that every flight adds $\frac{\delta t}{\Delta t}+1$ flight edges to the edge set, i.e.,
$\mathcal{A}^{f}_{t}=$${\Big{\{}}(i,j):i_{\mathrm{s}}=o_{f},\,j_{\mathrm{s}}=d_{f},\,i_{\mathrm{t}}=t_{f}^{\mathrm{sd}}+\tau\;\forall\tau\in\{0,\dots,\frac{\delta t}{\Delta t}\},$
$j_{\mathrm{t}}=i_{\mathrm{t}}+1{\Big{\}}}\quad\forall f:t_{f}^{\mathrm{sd}}\in\mathcal{N}\;.$
@@
$\forall\tau\in\{1,\ldots,t^{f}-1\}\Big{\}}\quad\forall(i,j)\in\mathcal{A}^{f}_{t},\,\forall f:t_{f}^{\mathrm{sd}}\in\mathcal{N}.$
@@ II-C Aircraft Routing
$[\mathbf{p}^{k}_{t}]_{i}=\begin{cases}-1\quad&i=\hat{p}_{\mathrm{M}}^{k}\\ 1\quad&i=z\\ 0\quad&\text{otherwise}\end{cases}\quad\forall k\in\mathcal{P}\;,$
@@
$[C_{t}]_{ab}=\begin{cases}1\quad&\{b=(i,j),\,a=(l,m):a\in\mathcal{C}_{b}\}\\ 0\quad&\text{otherwise}\end{cases}\;,$
@@
$[M_{t}]_{ab}=\begin{cases}\bar{M}\quad&a=(i,j)\in\mathcal{A}_{t},\,b=(i,j)\in\mathcal{A}_{t}^{\mathrm{g}}\\ 0\quad&\text{otherwise}\end{cases}\;,$
@@
$[A_{t}]_{fa}=\begin{cases}1\quad&a=(i,j)\in\mathcal{A}_{t}^{f}\\ 0\quad&\text{otherwise}\end{cases}\;.$
@@
$B_{t}\,\mathbf{x}^{k}=\mathbf{p}^{k}_{t}\quad$ $\forall k\in\mathcal{P}\;,$ (1)
$\mathbf{x}^{k}\geq C_{t}\,\mathbf{x}^{k}\quad$ $\forall k\in\mathcal{P}\;,$ (2)
$A_{t}\,\mathbf{x}=\mathbf{1}\quad$ if $\text{if }\exists f:t_{\mathrm{sd}}^{f}\in\mathcal{N}\;,$ (3)
$\mathbf{P}^{k}\leq M_{t}\,\mathbf{x}^{k}\quad$ $\forall k\in\mathcal{P}\;,$ (4)
$\mathbf{P}^{k}\leq M_{t}\,(\mathbf{1}-C_{t}\,\mathbf{x}^{k})\quad$ $\forall k\in\mathcal{P}\;,$ (5)
$[\mathbf{P}^{k}]_{i}\in[P_{\mathrm{c,min}},P_{\mathrm{c,max}}]\quad\forall i\in\mathcal{A}_{t}^{\mathrm{g}}\;$ $\forall k\in\mathcal{P}\;.$ (6)
@@
$[F_{t}]_{fa}=\begin{cases}\hat{E}_{f}\quad&a=(i,j)\in\mathcal{A}^{f}_{t}\\ 0\quad&\text{otherwise}\end{cases}\;.$
@@
$\mathbf{E}^{k}=E_{\mathrm{b,M}}^{k}+G_{t}\,\mathbf{P}^{k}\cdot\Delta t-F_{t}\,\mathbf{x}^{k}\quad$ $\forall k\in\mathcal{P}\;,$ (7)
$\mathbf{E}^{k}\geq[\mathbf{1}_{|\mathcal{H}|}^{\top}\otimes I_{N}]\,M_{t}\,C_{t}\,\mathbf{x}^{k}\cdot E_{\mathrm{res}}\quad$ $\forall k\in\mathcal{P}\;,$ (8)
$[\mathbf{E}^{k}]_{i}\in[E_{\mathrm{b,min}}^{k},E_{\mathrm{b,max}}^{k}]\quad\forall i\in\mathcal{N}\;$ $\forall k\in\mathcal{P}\;,$ (9)
$[\mathbf{E}^{k}]_{N}\in\mathcal{X}_{\mathbf{f}}^{E_{\mathrm{b}}}\quad$ $\forall k\in\mathcal{P}\;,$ (10)
@@ II-D Energy Model of the Airport
$\mathbf{E}_{\mathrm{b}}^{h}=E_{\mathrm{b,M}}^{h}-\Lambda_{N}\,\mathbf{P}_{\mathrm{b}}^{h}\cdot\Delta t\quad$ $\forall h\in\mathcal{H}\;,$ (11)
$\mathbf{P}_{\mathrm{gr}}^{h}\geq\mathbf{P}_{\mathrm{a}}^{h}-\hat{\mathbf{P}}_{\mathrm{rnw}}^{h}-\mathbf{P}_{\mathrm{b}}^{h}\quad$ $\forall h\in\mathcal{H}\;,$ (12)
$\mathbf{P}_{\mathrm{a}}^{h}=[\mathbf{1}_{|\mathcal{P}|}^{\top}\otimes I_{|\mathcal{H}|\cdot N}]\,\mathbf{P}_{\mathrm{c}}\quad$ $\forall h\in\mathcal{H}\;,$ (13)
$\mathbf{E}_{\mathrm{b}}^{h}\in[E_{\mathrm{b,min}}^{h},E_{\mathrm{b,max}}^{h}]\quad$ $\forall h\in\mathcal{H}\;.$ (14)
@@ II-E MPC Formulation
$[\mathbf{r}_{t}]_{a}=\begin{cases}\varphi\cdot(i_{\mathrm{t}}-t_{\mathrm{sd}}^{f})^{2}\quad&a=(i,j):(i,j)\in\mathcal{A}^{f}_{t}\\ 0\quad&\text{otherwise}\end{cases}\;,$
@@
Aircraft Assignment and Charge Scheduling The optimal aircraft assignment for the horizon is found through the solution of
@@
Figure 3: Scheme for implementation of the Aircraft Assignment and Charge Scheduling Problem. Blocks with rounded edges are exogenous inputs and variables refer to the respective quantity at the measured time unless otherwise specified.
III Numerical Simulation Results
We simulate our control framework on a commuter flight network at the Dutch ABC-Islands, Aruba, Bonaire, and Curaçao. It is currently served by a fleet of small aircraft that could be electrified in the future. With short flights and frequent demand, this scenario is a perfect proving ground for our control scheme. With real solar radiation data [24] and flight trajectories [25], we can simulate operations for a day in August, for which we have obtained the schedule from openly available data. We parse Problem 1 with YALMIP [26] and solve it with Gurobi [27] with a horizon of 2 hours and a discrete time step of 5 minutes, resulting in an average computation time of few seconds.
Figure 4: Example comparison of a current optimal input and state trajectories (solid lines) and the ones computed 5 time steps earlier (dashed lines) for the aircraft state of charge $\xi_{\mathrm{b}}^{k}$ and routing (top), and airport BESS state of charge $\xi_{\mathrm{b}}^{h}$ and powers (bottom). The vertical bar marks the current time step $t_{\mathrm{M}}$
@@ III-A Flight Simulations
We simulate the scheduled flights between airports using conventional flight profiles for other aircraft on the same routes courtesy of flightradar24.com [25]. We consider Eviation’s Alice [28] as the aircraft for our case study, parameters of which are estimated in [29]. We devise a quasi-static aircraft model, where first the required lift coefficient is obtained from the exogenous flight profile,
$c_{\mathrm{L}}(t)=\frac{2\cdot m}{\rho(t)\cdot S\cdot v^{2}(t)}\cdot(\dot{v}_{\mathrm{z}}(t)+g\cdot\cos\,\theta(t))\;,$ (15)
where $m$ and $S$ are the mass and wing area of the aircraft, $g$ is the gravitational acceleration, $\rho(t)$ is the air density at the current altitude, and $v(t)$, $v_{\mathrm{z}}(t)$, and $\theta(t)$ are the instantaneous velocity, its vertical component, and climb angle, respectively (cf. [30, C. 18]). The drag polar is modeled as a quadratic dependence on the lift, where the drag coefficient is
$c_{\mathrm{D}}(t)=c_{\mathrm{D,min}}+k\cdot c_{\mathrm{L}}^{2}(t)\;,$ (16)
with the constants $c_{\mathrm{D,min}}$ and $k$ estimated following a procedure from [30, C. 15]. Finally, the thrust required is derived from a force balance [30, C. 17-19]:
$F_{\mathrm{T}}(t)=m\cdot(v_{\mathrm{x}}(t)+g\cdot\sin\,\theta(t))+\frac{\rho(t)\cdot S\cdot c_{\mathrm{D}}(t)\cdot v^{2}(t)}{2}\;,$ (17)
where $v_{\mathrm{x}}(t)$ is the horizontal component of the velocity. We follow the procedure described in [30, C. 17] to find the take-off velocity and estimate the required thrust profile before take-off. With a constant power train efficiency as in [29], we find the energy required for the flight through
$E_{f}=\frac{1}{\eta_{\mathrm{pt}}}\cdot\int_{t_{\mathrm{d}}^{f}}^{t_{\mathrm{a}}^{f}}F_{\mathrm{T}}(t)\cdot v(t)\,\mathrm{d}t\;,$ (18)
where $t_{\mathrm{d}}^{f}$ and $t_{\mathrm{a}}^{f}$ are the departure and arrival time of flight $f$. Notice that whilst this model cannot be considered high-fidelity, as it neglects parameters such as the wind velocity, it still provides a good estimate of the energy consumption for a given flight trajectory [29].
@@ III-B Results
Results reported in Fig. 4 show the operation of the controller. Comparing the trajectories optimized at a current time step to those that were computed 5 time steps earlier, we find that aircraft re-assignment takes place as conditions change.
Rescheduling
In the example, the aircraft shown in orange arrives later than initially assumed and required more energy for its flight. As a consequence, its consecutive flight is delayed by one time step so it can be charged sufficiently. At the destination airport of that flight, Aruba, the blue aircraft is then charged while the orange one is en-route, such that it can serve a flight at 10:55, which was previously assigned to the orange aircraft that is instead kept at Aruba.
Energy Management
A deviation from the expected solar yield in Bonaire causes the yellow aircraft to be charged less than initially planned, instead it is now optimal to charge it at its destination airport later in the horizon. A previously planned significant dip in the BESS state of charge at Aruba at 10:45 is avoided by reassigning the aircraft that takes the flight at 10:55.
Discussion
A few comments are in order. First, the final sets for the battery state of charge of the aircraft is set very conservatively in this example, but further research on feasibility of the energy trajectories will address this. Second, reassigning aircraft at a moment’s notice may not always be practical, but the controller has shown that in cases where this cannot be avoided, it can successfully mitigate disturbances in the network.

### v2

@@ Abstract
This paper presents a Model Predictive Control (MPC) scheme for flight scheduling and energy management of electric aviation networks, where electric aircraft transport passengers between electrified airports equipped with sustainable energy sources and battery storage, with the goal of minimizing grid dependency. Specifically, we first model the aircraft flight and charge scheduling problem jointly with the airport energy management problem, explicitly accounting for local weather forecasts. Second, we frame the minimum-grid-energy operational problem as a mixed-integer linear program and solve it in a receding horizon fashion, where the route assignment and charging decisions of each aircraft can be dynamically reassigned to mitigate disruptions. We showcase the proposed MPC scheme on real-world data taken from a conventional flight network and weather conditions in the US American North East. The proposed framework saves between 10 and 37 % of grid energy requirements when compared to a baseline without re-routing. Hence, results show that MPC can effectively guarantee operation of the network by efficiently re-assigning flights and rescheduling aircraft charging, while maximizing the efficiency of the on-site energy systems.
@@ I Introduction
BESS $E_{\mathrm{b}}^{h}$ apron $E_{\mathrm{b}}^{1}$$E_{\mathrm{b}}^{k}$$P_{\mathrm{b}}$$P_{\mathrm{a}}$$P_{\mathrm{gr}}$$P_{\mathrm{rnw}}$$P_{\mathrm{c}}^{1}$$P_{\mathrm{c}}^{k}$
@@
In the following we often concatenate variables into column vectors denoted with bold letters as in $\mathbf{v}=[v(1)\;v(2)\;\dots\;v(n)]^{\top}$, such that the dimension of $\mathbf{v}$ would be $n\times 1$. A column vector of all ones with dimension $m$ is denoted $\mathbf{1}_{m}$, the $m\times m$ identity matrix is denoted $I_{m}$, and $\Lambda_{m}$ denotes an $m\times m$ lower triangular matrix of ones. Matrices are denoted with capital letters $M$, and the element in the $m$-th row and $n$-th column is indexed with $M]_{mn}$. We use $\otimes$ for the Kronecker product operator.
@@ II-B Aviation Network Model
We extract all flights that start in the upcoming horizon from the given flight schedule and construct a set of flight edges that connect nodes of different airports. For every flight we allow a maximum possible flight delay of $d_{\mathrm{max}}$, an integer multiple of $\Delta t$, meaning that every flight adds $\frac{d_{\mathrm{max}}}{\Delta t}+1$ flight edges to the edge set, i.e.,
$\mathcal{A}^{f}_{t}=$ ${\Big\{}(i,j):i_{\mathrm{s}}=o_{f},\,j_{\mathrm{s}}=d_{f},\,i_{\mathrm{t}}=t_{f}^{\mathrm{sd}}+\tau\;\forall\tau\in\{0,\dots,\frac{d_{\mathrm{max}}}{\Delta t}\},$
$j_{\mathrm{t}}=i_{\mathrm{t}}+1{\Big\}}\qquad\forall f:t_{f}^{\mathrm{sd}}\in\mathcal{N}\;.$
@@
$\forall\tau\in\{1,\ldots,t^{f}-1\}\Big\}\qquad\forall(i,j)\in\mathcal{A}^{f}_{t},\,\forall f:t_{f}^{\mathrm{sd}}\in\mathcal{N}.$
@@ II-C Aircraft Routing
$\mathbf{p}^{k}_{t}]_{i}=\begin{cases}-1\qquad&i=\hat{p}_{\mathrm{M}}^{k}\\ 1\qquad&i=z\\ 0\qquad&\text{otherwise}\end{cases}\quad\forall k\in\mathcal{P}\;,$
@@
$C_{t}]_{ab}=\begin{cases}1\qquad&\{b=(i,j),\,a=(l,m):a\in\mathcal{C}_{b}\}\\ 0\qquad&\text{otherwise}\end{cases}\;,$
@@
$M_{t}]_{ab}=\begin{cases}\bar{M}\qquad&a=(i,j)\in\mathcal{A}_{t},\,b=(i,j)\in\mathcal{A}_{t}^{\mathrm{g}}\\ 0\qquad&\text{otherwise}\end{cases}\;,$
@@
$A_{t}]_{fa}=\begin{cases}1\qquad&a=(i,j)\in\mathcal{A}_{t}^{f}\\ 0\qquad&\text{otherwise}\end{cases}\;.$
@@
$B_{t}\,\mathbf{x}^{k}=\mathbf{p}^{k}_{t}\qquad$ $\forall k\in\mathcal{P}\;,$ (1)
$\mathbf{x}^{k}\geq C_{t}\,\mathbf{x}^{k}\qquad$ $\forall k\in\mathcal{P}\;,$ (2)
$A_{t}\,\mathbf{x}=\mathbf{1}\qquad$ $\text{if }\exists f:t_{\mathrm{sd}}^{f}\in\mathcal{N}\;,$ (3)
$\mathbf{P}^{k}\leq M_{t}\,\mathbf{x}^{k}\qquad$ $\forall k\in\mathcal{P}\;,$ (4)
$\mathbf{P}^{k}\leq M_{t}\,(\mathbf{1}-C_{t}\,\mathbf{x}^{k})\qquad$ $\forall k\in\mathcal{P}\;,$ (5)
$[\mathbf{P}^{k}]_{i}\in[P_{\mathrm{c,min}},P_{\mathrm{c,max}}]\qquad\forall i\in\mathcal{A}_{t}^{\mathrm{g}}\;$ $\forall k\in\mathcal{P}\;.$ (6)
@@
$F_{t}]_{fa}=\begin{cases}\hat{E}_{f}\qquad&a=(i,j)\in\mathcal{A}^{f}_{t}\\ 0\qquad&\text{otherwise}\end{cases}\;.$
@@
$\mathbf{E}^{k}=E_{\mathrm{b,M}}^{k}+G_{t}\,\mathbf{P}^{k}\cdot\Delta t-F_{t}\,\mathbf{x}^{k}\qquad$ $\forall k\in\mathcal{P}\;,$ (7)
$\mathbf{E}^{k}\geq[\mathbf{1}_{|\mathcal{H}|}^{\top}\otimes I_{N}]\,M_{t}\,C_{t}\,\mathbf{x}^{k}\cdot E_{\mathrm{res}}\qquad$ $\forall k\in\mathcal{P}\;,$ (8)
$[\mathbf{E}^{k}]_{i}\in[E_{\mathrm{b,min}}^{k},E_{\mathrm{b,max}}^{k}]\qquad\forall i\in\mathcal{N}\;$ $\forall k\in\mathcal{P}\;,$ (9)
$[\mathbf{E}^{k}]_{N}\in\mathcal{X}_{\mathbf{f}}^{E_{\mathrm{b}}}\qquad$ $\forall k\in\mathcal{P}\;,$ (10)
@@ II-D Energy Model of the Airport
$\mathbf{E}_{\mathrm{b}}^{h}=E_{\mathrm{b,M}}^{h}-\Lambda_{N}\,\mathbf{P}_{\mathrm{b}}^{h}\cdot\Delta t\qquad$ $\forall h\in\mathcal{H}\;,$ (11)
$\mathbf{P}_{\mathrm{gr}}^{h}\geq\mathbf{P}_{\mathrm{a}}^{h}-\hat{\mathbf{P}}_{\mathrm{rnw}}^{h}-\mathbf{P}_{\mathrm{b}}^{h}\qquad$ $\forall h\in\mathcal{H}\;,$ (12)
$\mathbf{P}_{\mathrm{a}}^{h}=[\mathbf{1}_{|\mathcal{P}|}^{\top}\otimes I_{|\mathcal{H}|\cdot N}]\,\mathbf{P}_{\mathrm{c}}\qquad$ $\forall h\in\mathcal{H}\;,$ (13)
$\mathbf{E}_{\mathrm{b}}^{h}\in[E_{\mathrm{b,min}}^{h},E_{\mathrm{b,max}}^{h}]\qquad$ $\forall h\in\mathcal{H}\;.$ (14)
@@ II-E MPC Formulation
$\mathbf{r}_{t}]_{a}=\begin{cases}\varphi\cdot(i_{\mathrm{t}}-t_{\mathrm{sd}}^{f})^{2}\qquad&a=(i,j):(i,j)\in\mathcal{A}^{f}_{t}\\ 0\qquad&\text{otherwise}\end{cases}\;,$
@@
Aircraft Routing and Charge Scheduling (ARCS) The optimal aircraft assignment for the horizon is found through the solution of
@@
  ARCS       FleetAirportsAirportsAirports
solar
forecast
real solar power
build
graph
schedule
flight assignments,
$\{P^{k}\}_{k\in\mathcal{P}}$
$\{P_{\mathrm{b}}^{h},P_{\mathrm{gr}}^{h}\}_{h\in\mathcal{H}}$$\{P_{\mathrm{a}}^{h}\}_{h\in\mathcal{H}}$$\{P_{\mathrm{rnw}}^{h}\}_{h\in\mathcal{H}}$$\{P_{\mathrm{rnw}}^{h}\}_{h\in\mathcal{H}}$$\{\hat{P}_{\mathrm{rnw}}^{h}[t]\}_{h\in\mathcal{H},\atop t\in\mathcal{N}}$$\{f:t_{\mathrm{sd}}^{f}\in\mathcal{N}\}$$\mathcal{G}_{t},\mathrm{p}_{t}$$\{E_{\mathrm{b}}^{k},\hat{p}_{\mathrm{M}}^{k}\}_{k\in\mathrm{P}}$$\{E_{\mathrm{b}}^{h}\}_{h\in\mathcal{H}}$
Figure 3: Scheme for implementation of the Aircraft Routing and Charge Scheduling (ARCS). Blocks with rounded edges are exogenous inputs and variables refer to the respective quantity at the measured time unless otherwise specified.
III Numerical Simulations
We simulate our control framework on Cape Air’s commuter flight network in New England, where the airline operates around 200 daily flights between 14 airports as shown in Fig. 4. We use flights from a week in August 2024 [24] and obtain solar irradiation data in 10 min increments for every airport through the National Solar Radiation Data Base (NSRDB) [25].
We assume that every airport in the network has 1000 m2 of solar cells and a 500 kWh BESS. We parse Problem 1 with YALMIP [26] and solve it with Gurobi [27] with a horizon of 2 h and a discrete time step of 5 min, resulting in an average computation time of few seconds for $|\mathcal{A}_{t}|\cdot|\mathcal{P}|\approx 15,000$ binary variables per horizon.
Figure 4: Flight network operated by Cape Air in New England.
@@ III-A Flight Simulations
To simulate the scheduled flights between airports we adopt an agent-based approach with simplified quasi-static 2-D flight dynamics for every aircraft. Due to the absence of commercial electric aircraft at the time of writing, we envisage an electric version of the P2012 Traveler, inspired by [28] as the aircraft for our case study.
When a flight is assigned to an aircraft, a distance-based trajectory is computed that starts with a steady climb at a fixed angle to 1000 m, followed by a cruise phase at constant velocity and a descent once a certain way point before the destination airport is reached. The aircraft follows this trajectory as closely as possible, and every increment of time $\delta t$, the current thrust requirements are obtained through
$F_{\mathrm{t}}[k]=\frac{1}{2}\rho[k]S\left(c_{\mathrm{D,0}}+\frac{c_{\mathrm{L}}^{2}[k]}{\pi ARe}\right)v^{2}[k]+m\left(\sin(\theta[k])g+a[k]\right)\;,$
where $\rho[k]$ is the density at the current altitude, $S$ is the wing area, $c_{\mathrm{D,0}}$ the parasitic drag, $AR$ and $e$ the plane’s aspect ratio and Oswald efficiency, $\theta[k]$ the current climb angle, $v[k]$ and $a[k]$ are the velocity and acceleration in direction of travel, $m$ the mass and $g$ the gravitational constant [29, Ch’s. 17-22]. On the ground, an additional drag term is added, namely $(mg-\nicefrac{{1}}{{2}}\rho[k]Sc_{\mathrm{L}}[k]v^{2}[k])\mu$, where the friction coefficient $\mu$ includes braking during landing. The lift coefficient is
$c_{\mathrm{L}}[k]=\frac{2mg\cos(\theta[k])}{S\rho[k]v^{2}[k]}\;.$
We assume that during flight the airplane stays within its operational envelope, but restrict the deceleration during descent and landing such that the required lift coefficient does not exceed the maximum to prevent stalling. The propeller efficiency is interpolated from a look-up table
$\eta_{\mathrm{prop}}[k]=\mathcal{M}\left(P_{\mathrm{t}}[k],\rho[k],v[k]\right)\;,$
where $P_{\mathrm{t}}[k]=F_{\mathrm{t}}[k]v[k]$ is the thrust power (cf. [29, Ch. 14]). For the battery dynamics we assume a constant (dis-)charging efficiency, such that the state of energy can be obtained as
$E_{\mathrm{b}}[k+1]=\begin{cases}E_{\mathrm{b}}[k]-\frac{1}{\eta_{\mathrm{b}}\eta_{\mathrm{prop}}[k]}P_{\mathrm{t}}[k]\delta t\;&\text{during flight,}\\ E_{\mathrm{b}}[k]+\eta_{\mathrm{b}}P_{\mathrm{c}}[k]\delta t\;\;&\text{when charging.}\end{cases}$
We simulate the dynamics with $\delta t=1\,\text{s}$ and run the MPC whenever the clock time is a multiple of $\Delta t$.
@@ III-B Results
We show selected results for the case study at hand. First, we generate a baseline schedule by solving a fleet assignment problem similar to [11, 13] which we use to assign aircraft to their initial locations. We obtain grid power requirements for said baseline, where we do not allow for aircraft reassignment, but only energy management, which we optimize with our framework by enforcing its binary variables to follow the baseline schedule, although this baseline control scheme could be achieved with even simpler energy management schemes.
The proposed ARCS framework is applied as outlined in Fig. 3. Fig. 5 shows an example for a successful re-assignment: Initially, the plane flying flight 38 was then projected to be assigned to flight 58 after a period of recharging. However, the aircraft arrives earlier and at a lower SoC than expected, prompting a re-assignment of flight 58 to the yellow aircraft for which it is charged a little more and swaps flights with another aircraft that was already on the tarmac. Potentially, since flight 58 is a short flight to Boston (BOS), a higher SoC saves grid power there at a later time.
The BESS at the airport is used for charging the purple aircraft, which it was for the longest time not projected to do. This is likely caused by a change in solar irradiation conditions, and its rather unstable unstable prediction, which will be addressed in future work.
At airports with many aircraft present within a horizon, as is often the case at BOS, we find that frequent re-scheduling can take place, which may compromise passenger level of service. We plan to address this issue by exploring objective functions that penalize schedule deviations.
Figure 5: Comparison of projected (muted,dashed) and actual (solid) state of charge (SoC) trajectories for a selected time window at Martha’s Vineyard Airport (MVY). Flights 40/55 are from/to HPN, 40/(53,58) from/to BOS.
Figure 6: Comparison of grid energy requirements for electric aircraft operations at airports (with IATA identifiers) in the Cape Air flight network without (muted, background) and with (foreground) allowing for aircraft re-assignment.
Fig. 6 shows results for grid energy requirements for every day in the considered week compared to the baseline. Depending on schedule and weather conditions, we can save between 10 and 37 % of total grid energy in the whole network, thus making better use of the renewable energy sources installed at the airports, notably without causing flight delays. We find a large proportion of these savings at BOS, which is the airline’s main hub, with more than 80 % of their flights either starting or landing there. At other airports, grid energy requirements rise slightly with the proposed scheme in some cases, which may be owed to the objective function that merely penalizes the use of total grid power without taking local circumstances into account. These findings motivate investigation into objective functions using (forecast) energy costs [30] and sizing considerations for the energy infrastructure at specific airports.